### v1

# A Real-time Dyson Expansion Scheme: Efficient Inclusion of Dynamical Correlations in Non-equilibrium Spectral Properties

Abstract
Time resolved photoemission spectroscopy is the key technique to probe the real-time non-equilibrium dynamics of electronic states. Theoretical predictions of the time dependent spectral function for realistic systems is however, a challenge. Employing the Kadanoff-Baym equations to find this quantity results in a cubic scaling in the total number of time steps, quickly becoming prohibitive and often fail quantitatively and even qualitatively. In comparison, mean field methods have more favorable numerical scaling both in the number number of time steps and in the complexity associated with the cost of a evolving for a single time step, however they miss key spectral properties such as emergent spectral features. Here we present a scheme that allows for the inclusion of dynamical correlations to the spectral function while maintaining the same scaling in the number of time steps as for mean field approaches, while capturing the emergent physics. Further, the scheme can be efficiently implemented on top of equilibrium real-time many-body perturbation theory schemes and codes. We see excellent agreement with exact results for test systems. Furthermore we exemplify the method on a periodic system and demonstrate clear evidence that our proposed scheme produces complex spectral features including excitonic band replicas, features that are not observed using static mean field approaches.
preprint: APS/123-QED
Introduction
In recent years there has been great interest in the non-equilibrium properties of quantum materials. These systems can display phenomena such as ultrafast phase phase transitions[1, 2], formation of exotic states[3, 4] and allow tuning of material properties through driving[5, 6]. The first principles description of non-equilibrium dynamics of electrons, motivated by advances in experiments [7, 8, 9], has attracted attention in recent years [10, 11, 12] and stimulated new theoretical developments [13, 14, 15]. In particular the Hartree-Fock Generalized Kadanoff-Baym ansatz (HF-GKBA) emerged as a powerful tool as it offers an efficient way of computing certain properties non-equilibrium systems[16, 13, 10]. It has been shown in many scenarios to provide results that are as accurate as the full solution of the Kadanoff-Baym equations (KBEs)[17, 18, 19, 20, 21], which they approximate without the need to compute the Green’s function propagation on the full $t,t^{\prime}$ grid (illustrated in Fig 1)[13, 22]. Further, it does not suffer from the spurious damping that the KBEs sometimes display[23, 24]. There is however one major caveat to these observations: these excellent results are restricted to properties derived from the time diagonal component of the Green’s function. Since the HF-GKBA reconstructs the time off-diagonal components at the Hartree-Fock(HF) level, properties linked to these components perform similarly to HF results[25]. And, as is discussed in the supplemental information, in some cases the HF-GKBA results are even worse than HF alone[26].
Time off-diagonal components are however, the key ingredient in computing non-equilibrium time resolved spectral properties[27]. This function gives information about the non-equilibrium energy distribution of particles and holes in the system, and can be used observe phenomena such as exciton formation and non-equilibrium phase transitions[28, 29, 2]. Furthermore, to see the intra-gap electron-hole signatures it is necessary to include two body correlations when computing the time off-diagonals. A feature which we will show is missed by HF and the HF-GKBA.
Figure 1: Schematic of the off-diagonal GKBA scheme. In the two step procedure one first performs a time evolution along the time diagonal. The information from the diagonal is then used as a starting point to update the off-diagonal
While methods have been employed to improve the HF-GKBA, they have mainly been applied to improve the time diagonal and steady-state properties of the HF-GKBA[30, 31]. While there has been some application of these methods to improve spectral properties[32], the improvements do not go beyond the mean field level due to the underlying reconstruction of the off-diagonal. One obvious method to try improving these spectral features is to resort to the full KBE calculation[33, 34]. The downside of this is that the computational scaling is cubic in the number of time-steps, making it numerically impractical. Although methods exist to reduce the scaling of the KBEs[14, 35, 36, 37, 38], the KBEs can also suffer from artificial damping which leads to broadening of the spectral peaks and information loss. Here we overcome the numerical scaling issues of the KBEs by developing a new scheme based on a time off-diagonal expansion of the Green’s function that picks up features beyond the mean field level and demonstrates excellent agreement with exact results. Furthermore, our approach requires only a slight modification of any existing real-time many-body perturbation theory (MBPT) methodology [39, 40] and can thus leverage existing schemes to reduce the computational cost of Green’s function calculations.
In this letter we present a route to include dynamical correlations for the time off-diagonal Green’s function without the need to explicitly perform collision integrals. In fact, to compute the non-equilibrium spectral function one needs to evaluate the off-diagonal components of the Green’s function within a small neighborhood of the time-diagonal, centered around some time $t_{p}$, as shown schematically in Fig. 1. Here, the scaling for computing the non-equilibrium spectral function is $O(\tilde{N}_{t}^{2})$ , where $\tilde{N}_{t}$ is the number of time-points centered around the probe time $t_{p}$. In practice, $\tilde{N}_{t}$ is dependent on the windowing function, $\mathcal{S}(t)$
$\mathcal{S}(t-t_{p})=\frac{1}{\delta\sqrt{2\pi}}\mathrm{e}^{-\frac{(t-t_{p})^{2}}{2\delta^{2}}}$ (1)
which is related to the probe width and the time resolved spectral function:[27]
[…]
This windowing function determines the energy and temporal resolution of the non-equilibrium spectral function[41]. In practice $\tilde{N}_{t}\ll N_{t}$, leading to huge computational savings, which is constant and determined solely by the windowing function and the overall computational scaling is thus determined solely by the time-diagonal propagation.
Real-time Dyson expansion scheme
A simple proof included in the supplemental information shows the equivalence between the following two equations[26].
$\begin{split}&G^{\mathrm{R}}(12)=G^{\mathrm{R,MF}}(12)\\ &\hskip 42.67912pt+\int\int d3d4\medspace G^{\mathrm{R,MF}}(13)\Sigma^{\mathrm{R}}[G^{\mathrm{MF}}(34)]G^{\mathrm{R}}(42)\\ &\equiv\left[i\partial_{t_{1}}-h^{\mathrm{MF}}[G^{\mathrm{MF}}(1)]\right]G^{\mathrm{R}}(12)\\ &\hskip 85.35826pt-\int d3\medspace\Sigma^{\mathrm{R}}[G^{\mathrm{MF}}(13)]G^{\mathrm{R}}(32)=\delta(12)\end{split}$ (3)
The first equation, gives the Dyson equation describing a one-shot correction on-top of a reference (mean-field Hamiltonian) for an arbitrary self-energy, $\Sigma(t,t^{\prime})$[42]. The second is exactly the equation we will solve in this paper using a real-time Green’s function formalism. We note that at equilibrium our approach reduces to the standard MBPT correction on top of a mean field reference Hamiltonian. In practical calculations, such an approach typically leads to great improvement and even yields spectra containing signatures of strong dynamical correlations[43, 44, 40]. This equivalence builds an intuition behind the procedure: we use a mean-field method as an initial guess and we add a correction term that includes dynamical correlation effects through $\Sigma(t,t^{\prime})$. Our approach thus represents a natural extension of computational techniques applied in MBPT to non-equilibrium settings.
Using this form of the integro-differential equation for $G^{\mathrm{R}}(t,t^{\prime})$ and the equations of motion for $G^{\mathrm{HF}}(t,t^{\prime})$ we derive an ODE scheme that involves propagating $G^{\mathrm{R}}(t,t^{\prime})$ along with an auxiliary two-body propagator used to reconstruct the collision integral at each time step. A similar approach has also been applied to the time diagonal Green’s function in a linear scaling HF-GKBA scheme[13, 22]. Propagating the ODE scheme on the time off-diagonal for some $\tilde{N_{t}}$ steps scales as $O(\tilde{N}_{t})$. Thus reconstructing the off-diagonal from $\tilde{N}_{t}$ steps along the diagonal will indeed scale as $O(\tilde{N}^{2}_{t})$. Furthermore, the time evolution is independent for each starting point, making this scheme extremely easy to parallelize in time coordinates.
In the following, we only show the equations for $G^{\mathrm{R}}(t,t^{\prime})$ however we perform the same procedure for $G^{<}(t,t^{\prime})$. The details on this can be found in the supplemental information including the relevant equations and the derivations for both schemes[26]. From now on we will refer to this scheme as the real-time Dyson expansion (RT-DE).
The scheme involves solving the following equation for $G^{\mathrm{R}}(t,t^{\prime})$ along with a second ODE for a two-body propagator $\mathcal{F}_{ijkl}(t,t^{\prime})$.
$\begin{split}\frac{\mathrm{d}G^{\mathrm{R}}(t,t^{\prime})}{\mathrm{d}t}&=-i\left[h^{\mathrm{MF}}(t)G^{\mathrm{R}}(t,t^{\prime})+I^{\mathrm{R}}(t,t^{\prime})\right],\\ I^{\mathrm{R}}_{im}(t,t^{\prime})&=-\sum_{klp}w_{iklp}(t)\mathcal{F}_{lpmk}(t,t^{\prime}).\\ \end{split}$ (4)
Here, the retarded collision integral depends on the interaction tensor and a two-body propagator $\mathcal{F}_{ijkl}$. By employing the approximation in equation (3) we derive an ODE for $\mathcal{F}_{ijkl}$ given by
$\begin{split}\frac{\mathrm{d}\mathcal{F}_{lpmk}(t,t^{\prime})}{\mathrm{d}t}&=\sum_{qrsj}w_{qrsj}(t)\Big{[}G^{>}_{lq}(t)G^{<}_{pr}(t)G^{>}_{sk}(t)-G^{<}_{lq}(t)G^{<}_{pr}(t)G^{>}_{sk}(t)\Big{]}G^{\mathrm{R}}_{jm}(t,t^{\prime})-\\ &\hskip 142.26378pti\sum_{x}\Big{[}h^{\mathrm{MF}}_{lx}(t)\mathcal{F}_{xpmk}(t,t^{\prime})+h^{\mathrm{MF}}_{px}(t)\mathcal{F}_{lxmk}(t,t^{\prime})-\mathcal{F}_{lpmx}(t,t^{\prime})h^{\mathrm{MF}}_{xk}(t)\Big{]}.\\ \end{split}$ (5)
To fully demonstrate the the quality of this “one-shot correction” at equilibrium, we provide its numerical analysis in the supplemental information[26]. In this setup, the time diagonal terms of RT-DE represent a stationary solution of a mean-field Hamiltonian, while the time off-diagonals contain the information about the self-energy in the time domain. Hence, RT-DE is equivalent to the perturbative correction and we show that it is in excellent agreement with exact diagonalization. In contrast, the HF-GKBA spectral function is reconstructed by a static mean-field approach and, hence, practically identical to the results of the underlying mean-field Hamiltonian.
In the non-equilibrium setup the distinction between various methods becomes even more pronounced. Again, RT-DE requires a time evolution of $G^{<}(t,t)$ using a time-dependent mean-field method. Here, for simplicity, we use TD-HF on the time diagonal as a prototypical example. Note, however, that this scheme is applicable to any particular choice of the reference static Hamiltonian, including those beyond mean-field, from some form of a downfolding procedure. For completeness we also show in [26] that the difference between exact diagonalization and TD-HF is negligible on the diagonal. The time diagonal components are then inserted in equation (5). Starting from some $t^{\prime}=t$ we can time-step in the $t$ variable in the range $t^{\prime}<t<T_{\mathrm{max}}$. Here we again emphasize an important advantage of the RT-DE scheme. Since we can start from any point on the diagonal we choose to only compute those $\tilde{N}_{t}^{2}$ off-diagonal terms within the probe window.
Finally, we note that the time evolution require the following initial conditions (for each $t^{\prime}$), $G^{\mathrm{R}}(t^{\prime},t^{\prime})=-i$, $\mathcal{F}(t^{\prime},t^{\prime})=0$. For $G^{\mathrm{R}}(t,t^{\prime})$ the initial condition is exact and exactly represents the one-shot correction. However for the lesser component we need to select a starting point for $I^{<}(t,t)$ and $G^{<}(t,t)$ which, in principle, breaks the equivalence to the standard perturbative expansion. This matter is discussed in more detail in the supplemental information. For the results shown here we have taken $I^{<}(t,t)=0$ and $G^{<}(t,t)=G^{<,\mathrm{MF}}(t)$.
Application to non-equilibrium problems
We will now proceed by studying the non-equilibrium spectral function in equation (2) with our proposed scheme.
$\begin{split}\mathcal{H}&=-J\sum_{\langle i,j\rangle}c^{\dagger}_{i}c_{j}+U\sum_{i}n_{i\uparrow}n_{i\downarrow}+h^{\mathrm{N.E}}(t)\\ \end{split}$ (6)
First we apply the RT-DE to non-equilibrium system and compare to TD-HF and exact diagonalization results. We excite the system given by equation (6) with a short wavelength pulse given by $h_{ij}^{\mathrm{N.E}}(t)=\delta_{ij}E\cos\left(\frac{\pi r_{i}}{2}\right)\exp\left({-\frac{(t-t_{0})^{2}}{2T_{p}^{2}}}\right)$[17], where $r_{i}=\frac{N_{s}-1}{2}-i$ is the lattice vector for site $i$ and $N_{s}$ is the number of lattice sites.
Figure 2: Time resolved emission spectrum a) and full time dependent spectrum b) for TD-HF, exact diagonalization and the RT-DE. Calculated in the model given in equation 6 with the following parameters: $U=1.0J$, $N_{s}=4$, $E=0.5J$, $t_{0}=5J^{-1}$ and $T_{p}=0.5J^{-1}$ after time evolving to $T_{\mathrm{max}}=200J^{-1}$. The probe width is taken to be $\delta=10$ and $t_{p}=100J^{-1}$.
First we will examine Fig. 2 a). The RT-DE is in excellent agreement with the exact result for the lesser component of the spectral function. It picks up features around $\omega\approx-1.2J$ and $\omega\approx-0.4J$ that are completely missed by TD-HF. For the peaks above the Fermi energy, both the RT-DE and TD-HF pick up the peaks. Both methods slightly underestimate the energy of the peak around $\omega\approx 1.0J$ by the same amount. Interestingly, the peak near $\omega\approx 2.0J$ is capture slightly better with TD-HF than with the RT-DE. This is related to the initial condition for the lesser component. Given that the initial condition for $G^{<}(t,t^{\prime})$ can be fixed in different ways, further optimization and constraining the particular choices will be explored elsewhere. Yet, we emphasize that this boundary condition affects only the details of the electronic distribution in dynamically renormalized states. The band renormalization and emergence of new spectral peaks is however captured perfectly using the knowledge of the $G^{R}$ alone, which is free of any uncertainty with fixing its boundary conditions.
For the retarded component in Fig. 2 b), we see a similarly excellent of agreement between the exact result and the RT-DE. The RT-DE matches the non-equilibrium spectrum very closely, again it picks up peaks at $\omega\approx\pm 0.4J$ and $\omega\approx\pm 1.2J$ that are completely missed by HF and is in perfect agreement with exact results for the height of the peaks at $\omega\approx\pm 1.9J$. The mean-field method severely underestimates this spectral feature since it does not include dynamical effects that dominate the peak magnitude.
In our second example we explore a realistic model of an insulator out of equilibrium. We study the dynamics of a long-range interacting two band Hubbard model with periodic boundary conditions at half filling under a strong optical excitation. The full Hamiltonian is given by
$\begin{split}\mathcal{H}&=\sum_{\alpha,\beta\in\{c,v\}}\sum_{i,j,\sigma}h^{\alpha\beta}_{ij}(t)c^{\alpha\dagger}_{i,\sigma}c^{\beta}_{j,\sigma}+U\sum_{i}n^{\alpha}_{i\uparrow}n^{\alpha}_{i\uparrow}\\ &\hskip 71.13188pt+U\sum_{i}n_{ic}n_{iv}+\gamma U\sum_{i<j\atop\alpha\leq\beta}\frac{n_{i\alpha}n_{j\beta}}{|\vec{r}_{i}-\vec{r}_{j}|}.\\ h^{\alpha\beta}_{ij}(t)&=J\delta_{\alpha\beta}\delta_{\langle i,j\rangle}(\delta_{\alpha v}-\delta_{\alpha c})+\epsilon_{\alpha}\delta_{\alpha\beta}\\ &\hskip 42.67912pt+\delta_{ij}(1-\delta_{\alpha\beta})E\cos(\omega_{p}(t-t_{0}))\mathrm{e}^{-\frac{(t-t_{0})^{2}}{2T_{p}^{2}}}\end{split}$ (7)
Figure 3: Non-equilibrium time resolved emission spectrum for the model given in equation (7) with 12 sites evolved until time $T_{\mathrm{max}}=100J^{-1}$. Panels a), b) and c) show the spectra computed with TD-HF, the HF-GKBA and the RT-DE respectively. The model parameters are $U=4.0J$, $\gamma=0.5$, $\epsilon_{v}=-30J$, $\epsilon_{c}=-20J$, $E=2.5J$, $\omega_{p}=9.0J$, $t_{0}=10J^{-1}$, $T_{p}=1.8J^{-1}$. The probe width used in (2) is $\delta=5$ and the probe is centered at $t_{p}=90J^{-1}$. A prominent excitonic feature is seen for the RT-DE results just below the renormalized conduction band; further, a visible band satellite is visible at the bottom of the bandstructure.
The results are shown in Fig. 1. We compare TD-HF, the HF-GKBA and the RT-DE using the following model parameters: $U=4.0J$, $\epsilon_{v}=-30J$, $\epsilon_{c}=-20J$, $\gamma=0.5,E=2.5J$, $\omega_{p}=7.5J$, $t_{0}=10J^{-1}$ and $T_{p}=1.8J^{-1}$. The excitation frequency (pump), $\omega_{p}$ is below the equilibrium gap ($\sim 10.0J$) but it is chosen such that excitons can form in the system. We also remark that at equilibrium the three methods produce practically identical band structures, shown as red dashed lines in Fig. 3. This allows us to see how each method performs, purely from changes due to the system leaving equilibrium.
Fig. 3 a) shows the time resolved emission spectrum for TD-HF. The full non-equilibrium band structure is shown in blue dash-dotted lines. We note that due to the lack of dynamical self-energy (yielding possibly multiple quasiparticle solutions) no new (satellite) states can appear in the TD-HF result. The figure shows electrons are excited across the gap and into the conduction band around $k=0$. This change in the electon distribution leads to a renormalization of the band gap, which is governed by the changes in the orbital occupancies and related to changes in the time dependent density matrix. In practice, occupied states are shifted down in energy creating a relatively sharp discontinuity in the upper band which becomes partially filled around $k=0$. The concentration of electrons (holes) at the $k=0$ point of the conduction (valence) band and the mean field nature of TD-HF means the renormalization mainly effects the $k=0$ point, leading to a significant stretching of the bands. As expected the lack of two-body correlations in TD-HF means we observe no novel features in the spectral function which could be interpreted as emerging excitonic states beyond the band stretching.
In Fig. 3 b) we again show the time resolved emission spectrum, now computed using the HF-GKBA. This result shares features common to TD-HF as we expect because it is dominated by the off-diagonal terms dominated by HF. Firstly, while the HF signatures are diminished, we see a similar stretching of the bands relative to their equilibrium value. One reason for this change comes from an inconsistency between the treatment of the time diagonal and off-diagonal terms in HF-GKBA. Specifically it arises from the GKBA density matrix, $\rho^{\mathrm{GKBA}}(t)$, entering the HF equation for the off-diagonal instead of $\rho^{\mathrm{HF}}(t)$. When solving the TD-HF equation the density matrix will be projected on to a set of excited states of the HF Hamiltonian. Since in general at time $t$, $\rho^{\mathrm{GKBA}}(t)\neq\rho^{\mathrm{HF}}(t)$ the set of excited states we project onto will be different thus leading to a distinct energy spectrum. Such an inconsistency makes it a priori difficult to interpret HF-GKBA spectral results or to say how the difference from TD-HF will manifest. Another important observation is the lack of additional excitonic features. The picture is thus consistent with TD-HF (albeit with the magnitude of the dynamical band renormalization being significantly distinct). This is expected as the spectrum is dominated by the time off-diagonal terms treated at the HF level.
Finally we turn to Fig. 3 c), where we have computed the time resolved emission spectrum using the RT-DE scheme. The results differ markedly and show emergent features consistent with non-equilibrium dynamics. For the main renormalized bands, shown in blue dash-dotted lines, we see a picture somewhat similar to the HF-GKBA. Namely there is a renormalization of the band positions compared to equilibrium. Interestingly, for the RT-DE the renormalization effects both the $k\neq 0$ and $k=0$ points very similarly resulting in almost a rigid shift of the bands, compared to the stretching seen in Fig 3 a) and b). This is consistent with renormalized (increased) screening in the photoexcited system which affects the entire band structure and not only the small number states that become occupied by photo-excitation.
In addition, there are two more striking differences between the RT-DE and both TD-HF and the HF-GKBA. The first is a peak appearing just below the renormalized conduction band, consistent with an in-gap electronic state, i.e., an exciton. Secondly we see the appearance of a satellite band appearing just below the position of the equilibrium valence band. Similarly to the finite system in Fig. 2 the inclusion of dynamical correlations through the retarded collision integral allows for the formation of additional spectral features. This is consistent with satellites emerging due to coupling to a bosonic excitation in the system. Here, the satellite is displaced by energy equivalent to the distance between the excitonic feature and the minimum of the renormalized conduction state. The satellites thus can be interpreted as arising due to the simultaneous QP formation and exicton dissociation, i.e., corresponding to an “excitonic replica” of the valence band. The exact interpretation of such features are left for future research, but we note that these are consistent with the exact solutions made for finite systems studied above.
Conclusions and outlook
For our tests in finite systems the results for the RT-DE in and out of equilibrium are in excellent agreement with exact results. Not only does the RT-DE correct features picked up by TD-HF and the HF-GKBA but it also captures satellite features that both other methods miss entirely. For our larger system, the RT-DE results provide physically meaningful results. Importantly by including two-body correlations through the self-energy we are able to observe an emergent excitonic peak appearing in the spectral function and energetically well separated from the quasiparticle band.
In conclusion the RT-DE is an efficient scheme for studying the non-equilibrium spectral properties of quantum systems, offering accuracy and a huge reduction in cost relative to the full Kadanoff-Baym equations. Application of this scheme to a more realistic system and comparison of the results to experimental measurements as a validation is underway. Furthermore, deeper investigation into the new spectral features found for periodic system presented here is warranted. Finally, we note that RT-DE can be trivially combined with existing real-time MBPT methods and their extension to more advanced self-energy approximations, such as $GW\Gamma$[43].
Acknowledgements
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research and Office of Basic Energy Sciences, Scientific Discovery through Advanced Computing (SciDAC) program under Award Number DE-SC0022198. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC award BES-ERCAP0029462

### v2

@@
 Abstract
-Time resolved photoemission spectroscopy is the key technique to probe the real-time non-equilibrium dynamics of electronic states.
+Time-resolved photoemission spectroscopy is the key technique to probe the real-time non-equilibrium dynamics of electronic states.
 Theoretical predictions of the time dependent spectral function for realistic systems is however, a challenge.
 Employing the Kadanoff-Baym equations to find this quantity results in a cubic scaling in the total number of time steps, quickly becoming prohibitive and often fail quantitatively and even qualitatively.
-In comparison, mean field methods have more favorable numerical scaling both in the number number of time steps and in the complexity associated with the cost of a evolving for a single time step, however they miss key spectral properties such as emergent spectral features.
-Here we present a scheme that allows for the inclusion of dynamical correlations to the spectral function while maintaining the same scaling in the number of time steps as for mean field approaches, while capturing the emergent physics.
+In comparison, mean-field methods have more favorable numerical scaling both in the number of time steps and in the complexity associated with the cost of evolving for a single time step, however they miss key spectral properties such as emergent spectral features.
+Here we present a scheme that allows for the inclusion of dynamical correlations to the spectral function while maintaining the same scaling in the number of time steps as for mean-field approaches, while capturing the emergent physics.
 Further, the scheme can be efficiently implemented on top of equilibrium real-time many-body perturbation theory schemes and codes.
 We see excellent agreement with exact results for test systems.
-Furthermore we exemplify the method on a periodic system and demonstrate clear evidence that our proposed scheme produces complex spectral features including excitonic band replicas, features that are not observed using static mean field approaches.
+Furthermore we exemplify the method on a periodic system and demonstrate clear evidence that our proposed scheme produces complex spectral features including excitonic band replicas, features that are not observed using static mean-field approaches.
 preprint: APS/123-QED
@@
 In recent years there has been great interest in the non-equilibrium properties of quantum materials.
-These systems can display phenomena such as ultrafast phase phase transitions[1, 2], formation of exotic states[3, 4] and allow tuning of material properties through driving[5, 6].
+These systems can display phenomena such as ultrafast phase transitions[1, 2], formation of exotic states[3, 4] and allow tuning of material properties through driving[5, 6].
 The first principles description of non-equilibrium dynamics of electrons, motivated by advances in experiments [7, 8, 9], has attracted attention in recent years [10, 11, 12] and stimulated new theoretical developments [13, 14, 15].
-In particular the Hartree-Fock Generalized Kadanoff-Baym ansatz (HF-GKBA) emerged as a powerful tool as it offers an efficient way of computing certain properties non-equilibrium systems[16, 13, 10].
-It has been shown in many scenarios to provide results that are as accurate as the full solution of the Kadanoff-Baym equations (KBEs)[17, 18, 19, 20, 21], which they approximate without the need to compute the Green’s function propagation on the full $t,t^{\prime}$ grid (illustrated in Fig 1)[13, 22].
-Further, it does not suffer from the spurious damping that the KBEs sometimes display[23, 24].
-There is however one major caveat to these observations: these excellent results are restricted to properties derived from the time diagonal component of the Green’s function.
+In particular the Hartree-Fock Generalized Kadanoff-Baym ansatz (HF-GKBA) emerged as a powerful tool[16, 13, 10].
+It has been shown in many scenarios to provide results that are as accurate as the Kadanoff-Baym equations (KBEs)[17, 18, 19, 20, 21], which HF-GKBA approximates.
+Unlike the KBEs, the HF-GKBA does not require computation of the Green’s function (GF) on the full $t,t^{\prime}$ grid (Fig 1)[13, 22].
+Further, KBEs can suffer from spurious damping[23, 24].
+There is however one major caveat: these excellent results are restricted to properties derived from the time-diagonal component of the Green’s function.
 Since the HF-GKBA reconstructs the time off-diagonal components at the Hartree-Fock (HF) level, properties linked to these components perform similarly to HF results[25].
-And, as is discussed in the supplemental information, in some cases the HF-GKBA results are even worse than HF alone[26].
-Time off-diagonal components are however, the key ingredient in computing non-equilibrium time resolved spectral properties[27].
-This function gives information about the non-equilibrium energy distribution of particles and holes in the system, and can be used observe phenomena such as exciton formation and non-equilibrium phase transitions[28, 29, 2].
-Furthermore, to see the intra-gap electron-hole signatures it is necessary to include two body correlations when computing the time off-diagonals.
+In some cases the HF-GKBA results are even worse than HF alone[26].
+Time off-diagonal components are however, the key ingredient in computing the non-equilibrium time-resolved spectral function[27].
+This gives information about the non-equilibrium energy distribution of quasiparticles in the system, and captures exciton formation and non-equilibrium phase transitions[28, 29, 2].
+Furthermore, to see the intra-gap electron-hole signatures it is necessary to include two-body correlations when computing the time off-diagonals.
 A feature which we will show is missed by HF and the HF-GKBA.
 Figure 1: Schematic of the off-diagonal GKBA scheme.
-In the two step procedure one first performs a time evolution along the time diagonal.
+In the two step procedure one first performs a time evolution along the time-diagonal.
 The information from the diagonal is then used as a starting point to update the off-diagonal
-While methods have been employed to improve the HF-GKBA, they have mainly been applied to improve the time diagonal and steady-state properties of the HF-GKBA[30, 31].
-While there has been some application of these methods to improve spectral properties[32], the improvements do not go beyond the mean field level due to the underlying reconstruction of the off-diagonal.
-One obvious method to try improving these spectral features is to resort to the full KBE calculation[33, 34].
-The downside of this is that the computational scaling is cubic in the number of time-steps, making it numerically impractical.
-Although methods exist to reduce the scaling of the KBEs[14, 35, 36, 37, 38], the KBEs can also suffer from artificial damping which leads to broadening of the spectral peaks and information loss.
-Here we overcome the numerical scaling issues of the KBEs by developing a new scheme based on a time off-diagonal expansion of the Green’s function that picks up features beyond the mean field level and demonstrates excellent agreement with exact results.
-Furthermore, our approach requires only a slight modification of any existing real-time many-body perturbation theory (MBPT) methodology [39, 40] and can thus leverage existing schemes to reduce the computational cost of Green’s function calculations.
+While methods have been employed to improve the HF-GKBA, they have mainly been applied to improve the time-diagonal and steady-state properties[30, 31].
+Works that seek to improve the spectral properties[32] are scarce and the improvements do not go beyond the mean-field level due to the underlying reconstruction of the off-diagonal.
+Performing the full KBE calculations[33, 34] instead is however problematic as the computational scaling is cubic in the number of time-steps, making it numerically impractical.
+Although methods exist to reduce the scaling of the KBEs[14, 35, 36, 37, 38], they can also suffer from damping effects which leads to broadening of the spectral peaks and information loss.
+Here we overcome the numerical scaling issues of the KBEs by developing a new scheme that picks up features beyond the mean-field level, demonstrates excellent agreement with exact results and has effectively linear scaling in the number of time-steps.
 In this letter we present a route to include dynamical correlations for the time off-diagonal Green’s function without the need to explicitly perform collision integrals.
-In fact, to compute the non-equilibrium spectral function one needs to evaluate the off-diagonal components of the Green’s function within a small neighborhood of the time-diagonal, centered around some time $t_{p}$, as shown schematically in Fig.
+To compute the non-equilibrium spectral function one needs the off-diagonal components of the GF within a small neighborhood of the time-diagonal, centered around some time $t_{p}$, as shown schematically in Fig.
 1.
-Here, the scaling for computing the non-equilibrium spectral function is $O(\tilde{N}_{t}^{2})$ , where $\tilde{N}_{t}$ is the number of time-points centered around the probe time $t_{p}$.
-In practice, $\tilde{N}_{t}$ is dependent on the windowing function, $\mathcal{S}(t)$
-$\mathcal{S}(t-t_{p})=\frac{1}{\delta\sqrt{2\pi}}\mathrm{e}^{-\frac{(t-t_{p})^{2}}{2\delta^{2}}}$ (1)
-which is related to the probe width and the time resolved spectral function:[27]
+Here, the scaling for computing the non-equilibrium spectral function is at best $O(\tilde{N}_{t}^{2})$ , where $\tilde{N}_{t}$ is the number of time-points centered around the probe time $t_{p}$.
+In practice, $\tilde{N}_{t}$ is dependent on the windowing function, $\mathcal{S}(t)$ which is typically Gaussian and further discussed in the supplemental information (SI)[26].
+The time-resolved spectral function is expressed as,[27]
 […]
-This windowing function determines the energy and temporal resolution of the non-equilibrium spectral function[41].
-In practice $\tilde{N}_{t}\ll N_{t}$, leading to huge computational savings, which is constant and determined solely by the windowing function and the overall computational scaling is thus determined solely by the time-diagonal propagation.
-Real-time Dyson expansion scheme
-A simple proof included in the supplemental information shows the equivalence between the following two equations[26].
-$\begin{split}&G^{\mathrm{R}}(12)=G^{\mathrm{R,MF}}(12)\\ &\hskip 42.67912pt+\int\int d3d4\medspace G^{\mathrm{R,MF}}(13)\Sigma^{\mathrm{R}}[G^{\mathrm{MF}}(34)]G^{\mathrm{R}}(42)\\ &\equiv\left[i\partial_{t_{1}}-h^{\mathrm{MF}}[G^{\mathrm{MF}}(1)]\right]G^{\mathrm{R}}(12)\\ &\hskip 85.35826pt-\int d3\medspace\Sigma^{\mathrm{R}}[G^{\mathrm{MF}}(13)]G^{\mathrm{R}}(32)=\delta(12)\end{split}$ (3)
-The first equation, gives the Dyson equation describing a one-shot correction on-top of a reference (mean-field Hamiltonian) for an arbitrary self-energy, $\Sigma(t,t^{\prime})$[42].
-The second is exactly the equation we will solve in this paper using a real-time Green’s function formalism.
-We note that at equilibrium our approach reduces to the standard MBPT correction on top of a mean field reference Hamiltonian.
-In practical calculations, such an approach typically leads to great improvement and even yields spectra containing signatures of strong dynamical correlations[43, 44, 40].
-This equivalence builds an intuition behind the procedure: we use a mean-field method as an initial guess and we add a correction term that includes dynamical correlation effects through $\Sigma(t,t^{\prime})$.
-Our approach thus represents a natural extension of computational techniques applied in MBPT to non-equilibrium settings.
-Using this form of the integro-differential equation for $G^{\mathrm{R}}(t,t^{\prime})$ and the equations of motion for $G^{\mathrm{HF}}(t,t^{\prime})$ we derive an ODE scheme that involves propagating $G^{\mathrm{R}}(t,t^{\prime})$ along with an auxiliary two-body propagator used to reconstruct the collision integral at each time step.
-A similar approach has also been applied to the time diagonal Green’s function in a linear scaling HF-GKBA scheme[13, 22].
+Where $\mathcal{S}(t)$ determines the energy and temporal resolution of $\mathcal{A}(\omega,t_{p})$ [39].
+In practice $\tilde{N}_{t}\ll N_{t}$ and is constant and determined solely by $\mathcal{S}(t)$ and the overall computational scaling is thus determined solely by the time-diagonal propagation.
+The real-time Dyson expansion (RT-DE)
+In the SI we show that[26].
+$\begin{split}&G(12)=G^{\mathrm{MF}}(12)+\int\int d3d4\medspace G^{\mathrm{MF}}(13)\Sigma[G^{\mathrm{MF}}(34)]G(42)\\ &\equiv\left[i\partial_{t_{1}}-h^{\mathrm{MF}}[G^{\mathrm{MF}}(1)]\right]G(12)\\ &\hskip 71.13188pt-\int d3\medspace\Sigma[G^{\mathrm{MF}}(13)]G(32)=\delta(12)\end{split}$ (2)
+The first equation, gives the Dyson equation describing a one-shot correction on-top of a reference (arbitrary mean-field Hamiltonian (MFH)) for an arbitrary self-energy, $\Sigma(t,t^{\prime})$[40].
+The second is exactly the equation we will solve in this paper using a real-time GF formalism.
+At equilibrium, our approach reduces to the standard Many-body perturbation theory (MBPT) correction on top of a MFH.
+This typically leads to great improvement and even yields spectra containing signatures of strong dynamical correlations[41, 42, 43].
+Equation (2) builds an intuition behind the procedure: we use a mean-field method as an initial guess and we add a correction term that includes dynamical correlation effects through $\Sigma(t,t^{\prime})$.
+RT-DE thus represents a natural extension of the one-shot correction in MBPT to non-equilibrium settings.
+Using this form of the integro-differential equation for $G^{\mathrm{R}}(t,t^{\prime})$ and the equations of motion for $G^{\mathrm{MF}}(t,t^{\prime})$ we derive an ordinary differential equation (ODE) scheme that involves propagating $G^{\mathrm{R}}(t,t^{\prime})$ along with an auxiliary two-body propagator.
+A similar approach has also been applied to the time-diagonal GF in a linear scaling HF-GKBA scheme[13, 22].
 Propagating the ODE scheme on the time off-diagonal for some $\tilde{N_{t}}$ steps scales as $O(\tilde{N}_{t})$.
@@
 Furthermore, the time evolution is independent for each starting point, making this scheme extremely easy to parallelize in time coordinates.
-In the following, we only show the equations for $G^{\mathrm{R}}(t,t^{\prime})$ however we perform the same procedure for $G^{<}(t,t^{\prime})$.
-The details on this can be found in the supplemental information including the relevant equations and the derivations for both schemes[26].
-From now on we will refer to this scheme as the real-time Dyson expansion (RT-DE).
-The scheme involves solving the following equation for $G^{\mathrm{R}}(t,t^{\prime})$ along with a second ODE for a two-body propagator $\mathcal{F}_{ijkl}(t,t^{\prime})$.
-$\begin{split}\frac{\mathrm{d}G^{\mathrm{R}}(t,t^{\prime})}{\mathrm{d}t}&=-i\left[h^{\mathrm{MF}}(t)G^{\mathrm{R}}(t,t^{\prime})+I^{\mathrm{R}}(t,t^{\prime})\right],\\ I^{\mathrm{R}}_{im}(t,t^{\prime})&=-\sum_{klp}w_{iklp}(t)\mathcal{F}_{lpmk}(t,t^{\prime}).\\ \end{split}$ (4)
-Here, the retarded collision integral depends on the interaction tensor and a two-body propagator $\mathcal{F}_{ijkl}$.
-By employing the approximation in equation (3) we derive an ODE for $\mathcal{F}_{ijkl}$ given by
-$\begin{split}\frac{\mathrm{d}\mathcal{F}_{lpmk}(t,t^{\prime})}{\mathrm{d}t}&=\sum_{qrsj}w_{qrsj}(t)\Big{[}G^{>}_{lq}(t)G^{<}_{pr}(t)G^{>}_{sk}(t)-G^{<}_{lq}(t)G^{<}_{pr}(t)G^{>}_{sk}(t)\Big{]}G^{\mathrm{R}}_{jm}(t,t^{\prime})-\\ &\hskip 142.26378pti\sum_{x}\Big{[}h^{\mathrm{MF}}_{lx}(t)\mathcal{F}_{xpmk}(t,t^{\prime})+h^{\mathrm{MF}}_{px}(t)\mathcal{F}_{lxmk}(t,t^{\prime})-\mathcal{F}_{lpmx}(t,t^{\prime})h^{\mathrm{MF}}_{xk}(t)\Big{]}.\\ \end{split}$ (5)
-To fully demonstrate the the quality of this “one-shot correction” at equilibrium, we provide its numerical analysis in the supplemental information[26].
-In this setup, the time diagonal terms of RT-DE represent a stationary solution of a mean-field Hamiltonian, while the time off-diagonals contain the information about the self-energy in the time domain.
-Hence, RT-DE is equivalent to the perturbative correction and we show that it is in excellent agreement with exact diagonalization.
-In contrast, the HF-GKBA spectral function is reconstructed by a static mean-field approach and, hence, practically identical to the results of the underlying mean-field Hamiltonian.
-In the non-equilibrium setup the distinction between various methods becomes even more pronounced.
-Again, RT-DE requires a time evolution of $G^{<}(t,t)$ using a time-dependent mean-field method.
-Here, for simplicity, we use TD-HF on the time diagonal as a prototypical example.
-Note, however, that this scheme is applicable to any particular choice of the reference static Hamiltonian, including those beyond mean-field, from some form of a downfolding procedure.
-For completeness we also show in [26] that the difference between exact diagonalization and TD-HF is negligible on the diagonal.
-The time diagonal components are then inserted in equation (5).
-Starting from some $t^{\prime}=t$ we can time-step in the $t$ variable in the range $t^{\prime}<t<T_{\mathrm{max}}$.
+In the following, we only show the equations for $G^{\mathrm{<}}(t,t^{\prime})$ however we perform the same procedure for $G^{\mathrm{R}}(t,t^{\prime})$.
+The details are in the SI, which includes the derivation of the RT-DE scheme[26].
+The RT-DE involves solving the following equation for $G^{<}(t,t^{\prime})$.
+$\begin{split}\frac{\mathrm{d}G^{\mathrm{<}}(t,t^{\prime})}{\mathrm{d}t}&=-i\left[h^{\mathrm{MF}}(t)G^{\mathrm{<}}(t,t^{\prime})+I^{\mathrm{<}}(t,t^{\prime})\right],\\ I^{\mathrm{<}}_{im}(t,t^{\prime})&=-\sum_{klp}w_{iklp}(t)\mathcal{F}_{lpmk}(t,t^{\prime}).\\ \end{split}$ (3)
+Here, the lesser collision integral, $I^{<}(t,t^{\prime})$, depends on the interaction tensor and a two-body propagator $\mathcal{F}_{ijkl}$.
+By employing the approximation in equation (2) we introduce an ODE for $\mathcal{F}_{ijkl}$ given by
+$\begin{split}\frac{\mathrm{d}\mathcal{F}_{lpmk}(t,t^{\prime})}{\mathrm{d}t}&=\sum_{qrsj}w_{qrsj}(t)\Big{[}G^{>,\textrm{MF}}_{lq}(t)G^{<,\textrm{MF}}_{pr}(t)G^{>,\textrm{MF}}_{sk}(t)-G^{<,\textrm{MF}}_{lq}(t)G^{<,\textrm{MF}}_{pr}(t)G^{<,\textrm{MF}}_{sk}(t)\Big{]}G^{\mathrm{<}}_{jm}(t,t^{\prime})\\ &\hskip 142.26378pt-i\sum_{x}\Big{[}h^{\mathrm{MF}}_{lx}(t)\mathcal{F}_{xpmk}(t,t^{\prime})+h^{\mathrm{MF}}_{px}(t)\mathcal{F}_{lxmk}(t,t^{\prime})-\mathcal{F}_{lpmx}(t,t^{\prime})h^{\mathrm{MF}}_{xk}(t)\Big{]}.\\ \end{split}$ (4)
+To fully demonstrate the the quality of this “one-shot correction” at equilibrium, we provide its numerical analysis in the SI, where we see excellent results[26].
+Additionally we discuss the failings of mean-field approaches like HF and the HF-GKBA
+In the non-equilibrium setting the RT-DE requires a time evolution of $G^{<}(t,t)$ using a time-dependent MFH.
+While our scheme is applicable to any particular choice of the reference static Hamiltonian, such as from a time-dependent density functional theory calculation[44] and even those beyond mean-field e.g from some form of a downfolding procedure, we apply TD-HF on the time-diagonals for simplicity.
+This is discussed further in the SI[26].
+For completeness, we also show in [26] that the difference between exact diagonalization and TD-HF is relatively small on the diagonal.
+The time-diagonal components are inserted in equation (4).
+Starting from some $t^{\prime}=t$ we can time-step in the $t$ variable in the range $t^{\prime}\leq t\leq T_{\mathrm{max}}$.
 Here we again emphasize an important advantage of the RT-DE scheme.
 Since we can start from any point on the diagonal we choose to only compute those $\tilde{N}_{t}^{2}$ off-diagonal terms within the probe window.
-Finally, we note that the time evolution require the following initial conditions (for each $t^{\prime}$), $G^{\mathrm{R}}(t^{\prime},t^{\prime})=-i$, $\mathcal{F}(t^{\prime},t^{\prime})=0$.
-For $G^{\mathrm{R}}(t,t^{\prime})$ the initial condition is exact and exactly represents the one-shot correction.
-However for the lesser component we need to select a starting point for $I^{<}(t,t)$ and $G^{<}(t,t)$ which, in principle, breaks the equivalence to the standard perturbative expansion.
-This matter is discussed in more detail in the supplemental information.
+Finally, we we need to select an initial condition for $I^{<}(t,t)$ and $G^{<}(t,t)$ which, in principle, can break the equivalence to the standard perturbative expansion.
+This matter is discussed in more detail in the SI, as well as the initial conditions for the retarded component[26].
 For the results shown here we have taken $I^{<}(t,t)=0$ and $G^{<}(t,t)=G^{<,\mathrm{MF}}(t)$.
 Application to non-equilibrium problems
-We will now proceed by studying the non-equilibrium spectral function in equation (2) with our proposed scheme.
-$\begin{split}\mathcal{H}&=-J\sum_{\langle i,j\rangle}c^{\dagger}_{i}c_{j}+U\sum_{i}n_{i\uparrow}n_{i\downarrow}+h^{\mathrm{N.E}}(t)\\ \end{split}$ (6)
-First we apply the RT-DE to non-equilibrium system and compare to TD-HF and exact diagonalization results.
+We will now proceed by studying $\mathcal{A}(\omega,t_{p})$ with our proposed scheme.
+In this manuscript we only discuss comparisons of $\mathcal{A}(\omega,t_{p})$ since this is the experimentally relevant quantity extracted from the time off-diagonal GF.
+Further, this serves to compress the information stored in the GF making our comparisons clearer.
+$\begin{split}\mathcal{H}&=-J\sum_{\langle i,j\rangle}c^{\dagger}_{i}c_{j}+U\sum_{i}n_{i\uparrow}n_{i\downarrow}+\sum_{ij}h_{ij}^{\mathrm{N.E}}(t)c_{i}^{\dagger}c_{j}\\ \end{split}$ (5)
+First, to demonstrate the accuracy of our method, we apply the RT-DE to non-equilibrium system and compare to TD-HF and exact diagonalization.
 We excite the system given by equation (5) with a short wavelength pulse given by $h_{ij}^{\mathrm{N.E}}(t)=\delta_{ij}E\cos\left(\frac{\pi r_{i}}{2}\right)\exp\left({-\frac{(t-t_{0})^{2}}{2T_{p}^{2}}}\right)$[17], where $r_{i}=\frac{N_{s}-1}{2}-i$ is the lattice vector for site $i$ and $N_{s}$ is the number of lattice sites.
-Figure 2: Time resolved emission spectrum a) and full time dependent spectrum b) for TD-HF, exact diagonalization and the RT-DE.
-Calculated in the model given in equation 6 with the following parameters: $U=1.0J$, $N_{s}=4$, $E=0.5J$, $t_{0}=5J^{-1}$ and $T_{p}=0.5J^{-1}$ after time evolving to $T_{\mathrm{max}}=200J^{-1}$.
+Results for the lesser and retarded spectrum for dynamics induced by a long wavelength electric field and by a quench of the system are discussed in the SI[26].
+Figure 2: Time-resolved emission spectrum exact diagonalization (ED), RT-DE which is in excellent agreement with ED, and HF-GKBA that mostly copies the TD Hartree Fock (TD-HF) spectrum.
+Calculated in the model given in equation (5) with the following parameters: $U=1.0J$, $N_{s}=4$, $E=0.5J$, $t_{0}=5J^{-1}$ and $T_{p}=0.5J^{-1}$ after time evolving to $T_{\mathrm{max}}=200J^{-1}$.
 The probe width is taken to be $\delta=10$ and $t_{p}=100J^{-1}$.
-First we will examine Fig.
-2 a).
-The RT-DE is in excellent agreement with the exact result for the lesser component of the spectral function.
-It picks up features around $\omega\approx-1.2J$ and $\omega\approx-0.4J$ that are completely missed by TD-HF.
-For the peaks above the Fermi energy, both the RT-DE and TD-HF pick up the peaks.
-Both methods slightly underestimate the energy of the peak around $\omega\approx 1.0J$ by the same amount.
+Examining Fig.
+2 we see the RT-DE is in excellent agreement with the exact result for the lesser component of the $\mathcal{A}(\omega,t_{p})$.
+It picks up features around $\omega\approx-1.2J$ and $\omega\approx-0.4J$ that are completely missed by TD-HF and HF-GKBA, which mostly copies the TD-HF features.
+Notably, neither of these peaks is present in equilibrium[26].
+This shows the RT-DE can pick up emergent features that arise purely due to non-equilibrium effects in the system.
+For the peaks above the Fermi energy,the RT-DE, HF-GKBA and TD-HF pick up the peaks at $\omega\approx 1.0J$ and $2.0J$.
+TD-HF and the RT-DE methods slightly underestimate the energy of the peak around $\omega\approx 1.0J$ by the same amount.
 Interestingly, the peak near $\omega\approx 2.0J$ is capture slightly better with TD-HF than with the RT-DE.
 This is related to the initial condition for the lesser component.
-Given that the initial condition for $G^{<}(t,t^{\prime})$ can be fixed in different ways, further optimization and constraining the particular choices will be explored elsewhere.
-Yet, we emphasize that this boundary condition affects only the details of the electronic distribution in dynamically renormalized states.
-The band renormalization and emergence of new spectral peaks is however captured perfectly using the knowledge of the $G^{R}$ alone, which is free of any uncertainty with fixing its boundary conditions.
-For the retarded component in Fig.
-2 b), we see a similarly excellent of agreement between the exact result and the RT-DE.
-The RT-DE matches the non-equilibrium spectrum very closely, again it picks up peaks at $\omega\approx\pm 0.4J$ and $\omega\approx\pm 1.2J$ that are completely missed by HF and is in perfect agreement with exact results for the height of the peaks at $\omega\approx\pm 1.9J$.
-The mean-field method severely underestimates this spectral feature since it does not include dynamical effects that dominate the peak magnitude.
+Given that the initial condition for $G^{<}(t,t^{\prime})$ can be fixed in different ways.
+Further optimization and constraining the particular choices will be explored elsewhere.
+However, this boundary condition affects only the details of the electronic distribution in dynamically renormalized states.
+The band renormalization and emergence of new spectral peaks is however captured perfectly using the knowledge of the $G^{R}$ alone, which is free of any uncertainty.
+Finally, we note the appearance of two additional peaks in the HF-GKBA result at $\omega\approx 0.5J$ and $\omega\approx 1.6J$
+For completeness we provide discussion of the full spectrum from $G^{\mathrm{R}}(t,t^{\prime})$ in the SI[26].
 In our second example we explore a realistic model of an insulator out of equilibrium.
@@
 The full Hamiltonian is given by
-$\begin{split}\mathcal{H}&=\sum_{\alpha,\beta\in\{c,v\}}\sum_{i,j,\sigma}h^{\alpha\beta}_{ij}(t)c^{\alpha\dagger}_{i,\sigma}c^{\beta}_{j,\sigma}+U\sum_{i}n^{\alpha}_{i\uparrow}n^{\alpha}_{i\uparrow}\\ &\hskip 71.13188pt+U\sum_{i}n_{ic}n_{iv}+\gamma U\sum_{i<j\atop\alpha\leq\beta}\frac{n_{i\alpha}n_{j\beta}}{|\vec{r}_{i}-\vec{r}_{j}|}.\\ h^{\alpha\beta}_{ij}(t)&=J\delta_{\alpha\beta}\delta_{\langle i,j\rangle}(\delta_{\alpha v}-\delta_{\alpha c})+\epsilon_{\alpha}\delta_{\alpha\beta}\\ &\hskip 42.67912pt+\delta_{ij}(1-\delta_{\alpha\beta})E\cos(\omega_{p}(t-t_{0}))\mathrm{e}^{-\frac{(t-t_{0})^{2}}{2T_{p}^{2}}}\end{split}$ (7)
-Figure 3: Non-equilibrium time resolved emission spectrum for the model given in equation (7) with 12 sites evolved until time $T_{\mathrm{max}}=100J^{-1}$.
+$\begin{split}\mathcal{H}&=\sum_{\alpha,\beta\in\{c,v\}}\sum_{\langle i,j\rangle,\sigma}h^{\alpha\beta}_{ij}(t)c^{\dagger\alpha}_{i,\sigma}c^{\beta}_{j,\sigma}+U\sum_{i,\alpha}n^{\alpha}_{i\uparrow}n^{\alpha}_{i\downarrow}\\ &\hskip 71.13188pt+U\sum_{i}n_{ic}n_{iv}+\gamma U\sum_{i<j\atop\alpha\leq\beta}\frac{n_{i\alpha}n_{j\beta}}{|\vec{r}_{i}-\vec{r}_{j}|}.\\ h^{\alpha\beta}_{ij}(t)&=J\delta_{\alpha\beta}\delta_{\langle i,j\rangle}(\delta_{\alpha v}-\delta_{\alpha c})+\epsilon_{\alpha}\delta_{\alpha\beta}\\ &\hskip 42.67912pt+\delta_{ij}(1-\delta_{\alpha\beta})E\cos(\omega_{p}(t-t_{0}))\mathrm{e}^{-\frac{(t-t_{0})^{2}}{2T_{p}^{2}}}\end{split}$ (6)
+Figure 3: Non-equilibrium time-resolved emission spectrum for the model given in equation (6) with 12 sites evolved until time $T_{\mathrm{max}}=100J^{-1}$.
 Panels a), b) and c) show the spectra computed with TD-HF, the HF-GKBA and the RT-DE respectively.
-The model parameters are $U=4.0J$, $\gamma=0.5$, $\epsilon_{v}=-30J$, $\epsilon_{c}=-20J$, $E=2.5J$, $\omega_{p}=9.0J$, $t_{0}=10J^{-1}$, $T_{p}=1.8J^{-1}$.
+The model parameters are $U=4.0J$, $\gamma=0.5$, $\epsilon_{v}=-30J$, $\epsilon_{c}=-20J$, $E=2.5J$, $\omega_{p}=7.5J$, $t_{0}=10J^{-1}$, $T_{p}=1.8J^{-1}$.
 The probe width used in (1) is $\delta=5$ and the probe is centered at $t_{p}=90J^{-1}$.
@@
 The results are shown in Fig.
-1.
+3.
 We compare TD-HF, the HF-GKBA and the RT-DE using the following model parameters: $U=4.0J$, $\epsilon_{v}=-30J$, $\epsilon_{c}=-20J$, $\gamma=0.5,E=2.5J$, $\omega_{p}=7.5J$, $t_{0}=10J^{-1}$ and $T_{p}=1.8J^{-1}$.
@@
 This allows us to see how each method performs, purely from changes due to the system leaving equilibrium.
+The periodic system is too large for benchmarking against exact results; hence, this example serves primarily as a demonstration of the ability of the RT-DE to be applied to relatively large periodic systems.
+Even in the absence of the exact reference, we compare the qualitative features of $\mathcal{A}(\omega,t_{p})$ and determine the physical origins of the dynamical features that are produced by RT-DE (but not by the other methods).
 Fig.
-3 a) shows the time resolved emission spectrum for TD-HF.
+3 a) shows $\mathcal{A}(\omega,t_{p})$ for TD-HF.
 The full non-equilibrium band structure is shown in blue dash-dotted lines.
-We note that due to the lack of dynamical self-energy (yielding possibly multiple quasiparticle solutions) no new (satellite) states can appear in the TD-HF result.
+Due to the lack of dynamical self-energy (yielding possibly multiple quasiparticle solutions) no new (satellite) states can appear in the TD-HF result.
 The figure shows electrons are excited across the gap and into the conduction band around $k=0$.
-This change in the electon distribution leads to a renormalization of the band gap, which is governed by the changes in the orbital occupancies and related to changes in the time dependent density matrix.
+This change in the electron distribution leads to a renormalization of the band gap, which is governed by the changes in the orbital occupancies and related to changes in the time dependent density matrix.
 In practice, occupied states are shifted down in energy creating a relatively sharp discontinuity in the upper band which becomes partially filled around $k=0$.
-The concentration of electrons (holes) at the $k=0$ point of the conduction (valence) band and the mean field nature of TD-HF means the renormalization mainly effects the $k=0$ point, leading to a significant stretching of the bands.
-As expected the lack of two-body correlations in TD-HF means we observe no novel features in the spectral function which could be interpreted as emerging excitonic states beyond the band stretching.
+The concentration of electrons (holes) at the the conduction (valence) band extrema and the mean-field nature of TD-HF means the renormalization mainly effects the $k=0$ point, leading to a significant stretching of the bands.
+As expected the lack of two-body correlations in TD-HF means we observe no novel features in $\mathcal{A}(\omega,t_{p})$ which could be interpreted as emerging excitonic states beyond the band stretching.
 In Fig.
-3 b) we again show the time resolved emission spectrum, now computed using the HF-GKBA.
+3 b) we again show $\mathcal{A}(\omega,t_{p})$, now computed using the HF-GKBA.
 This result shares features common to TD-HF as we expect because it is dominated by the off-diagonal terms dominated by HF.
 Firstly, while the HF signatures are diminished, we see a similar stretching of the bands relative to their equilibrium value.
-One reason for this change comes from an inconsistency between the treatment of the time diagonal and off-diagonal terms in HF-GKBA.
-Specifically it arises from the GKBA density matrix, $\rho^{\mathrm{GKBA}}(t)$, entering the HF equation for the off-diagonal instead of $\rho^{\mathrm{HF}}(t)$.
-When solving the TD-HF equation the density matrix will be projected on to a set of excited states of the HF Hamiltonian.
-Since in general at time $t$, $\rho^{\mathrm{GKBA}}(t)\neq\rho^{\mathrm{HF}}(t)$ the set of excited states we project onto will be different thus leading to a distinct energy spectrum.
-Such an inconsistency makes it a priori difficult to interpret HF-GKBA spectral results or to say how the difference from TD-HF will manifest.
-Another important observation is the lack of additional excitonic features.
-The picture is thus consistent with TD-HF (albeit with the magnitude of the dynamical band renormalization being significantly distinct).
-This is expected as the spectrum is dominated by the time off-diagonal terms treated at the HF level.
+One reason for this change comes from an inconsistency between the treatment of the time-diagonal and off-diagonal terms in HF-GKBA, as the HF-GKBA density matrix, $\rho^{\mathrm{HF-GKBA}}(t)$, enters the HF equation for the off-diagonal instead of $\rho^{\mathrm{HF}}(t)$.
+The density matrix is thus projected on to a set of excited states of the HF Hamiltonian.
+Such an inconsistency makes it a priori difficult to interpret HF-GKBA spectral results.
+Importantly the HF-GKBA again lacks additional excitonic features.
+The picture is thus consistent with TD-HF (albeit with distinct band renormalization ).
 Finally we turn to Fig.
-3 c), where we have computed the time resolved emission spectrum using the RT-DE scheme.
-The results differ markedly and show emergent features consistent with non-equilibrium dynamics.
-For the main renormalized bands, shown in blue dash-dotted lines, we see a picture somewhat similar to the HF-GKBA.
-Namely there is a renormalization of the band positions compared to equilibrium.
-Interestingly, for the RT-DE the renormalization effects both the $k\neq 0$ and $k=0$ points very similarly resulting in almost a rigid shift of the bands, compared to the stretching seen in Fig 3 a) and b).
-This is consistent with renormalized (increased) screening in the photoexcited system which affects the entire band structure and not only the small number states that become occupied by photo-excitation.
-In addition, there are two more striking differences between the RT-DE and both TD-HF and the HF-GKBA.
+3 c), showing the RT-DE results.
+The results differ markedly and show emergent features.
+For the main renormalized bands we see a renormalization of the band positions compared to equilibrium.
+This affects both the $k\neq 0$ and $k=0$ points very similarly resulting in almost a rigid shift of the bands, compared to the stretching seen in Fig 3 a) and b).
+This is consistent with renormalized (increased) screening in the photoexcited system affecting the entire band structure and not only the small number states that become occupied by photo-excitation.
+In addition, there are two more striking differences.
 The first is a peak appearing just below the renormalized conduction band, consistent with an in-gap electronic state, i.e., an exciton.
@@
 Similarly to the finite system in Fig.
-2 the inclusion of dynamical correlations through the retarded collision integral allows for the formation of additional spectral features.
+2 the inclusion of dynamical correlations through $I^{<}(t,t^{\prime})$ allows for the formation of additional spectral features.
 This is consistent with satellites emerging due to coupling to a bosonic excitation in the system.
 Here, the satellite is displaced by energy equivalent to the distance between the excitonic feature and the minimum of the renormalized conduction state.
-The satellites thus can be interpreted as arising due to the simultaneous QP formation and exicton dissociation, i.e., corresponding to an “excitonic replica” of the valence band.
+Potentially, one can interpret this as an “excitonic replica”, i.e., simulaneous QP formation and exciton dissociation, similar to other bosonic replicas[42].
 The exact interpretation of such features are left for future research, but we note that these are consistent with the exact solutions made for finite systems studied above.
@@
 Not only does the RT-DE correct features picked up by TD-HF and the HF-GKBA but it also captures satellite features that both other methods miss entirely.
-For our larger system, the RT-DE results provide physically meaningful results.
-Importantly by including two-body correlations through the self-energy we are able to observe an emergent excitonic peak appearing in the spectral function and energetically well separated from the quasiparticle band.
-In conclusion the RT-DE is an efficient scheme for studying the non-equilibrium spectral properties of quantum systems, offering accuracy and a huge reduction in cost relative to the full Kadanoff-Baym equations.
+It is important to highlight that these excellent results are produced using the second Born approximation (representing one of the simplest self-energy forms).
+Thus, the results and comparisons to exact diagonalization given in this manuscript serve as compelling a proof of principle of the RT-DE.
+Already our method can capture dynamical correlation effects such as plasmons to a high degree of accuracy and it is natural to expect that extending to more advanced self-energy approximations will only further improve upon the results presented here, an area that is currently being pursued.
+To show the generality of our approach we have provided a derivation of the RT-DE for the $GW$ self-energy in the SI[26], it’s implementation, testing and comparison is left to a later publication.
+For our periodic system, the RT-DE provides physically meaningful results.
+Importantly by including two-body correlations through $\Sigma(t,t^{\prime})$ we are able to observe an emergent excitonic peak appearing in $\mathcal{A}(\omega,t_{p})$ and energetically well separated from the quasiparticle band.
+In conclusion the RT-DE is an efficient scheme for studying the non-equilibrium spectral properties of quantum systems, offering accuracy and a huge reduction in cost relative to the full KBEs.
+While KBEs and RT-DE will scale similarly in system size, by reducing the cost from cubic to linear in the number of time-steps, we overcome one of the major limitations of the KBEs that has prevented them being widely applied to non-equilibrium first principles simulations.
 Application of this scheme to a more realistic system and comparison of the results to experimental measurements as a validation is underway.
-Furthermore, deeper investigation into the new spectral features found for periodic system presented here is warranted.
-Finally, we note that RT-DE can be trivially combined with existing real-time MBPT methods and their extension to more advanced self-energy approximations, such as $GW\Gamma$[43].
+Finally, we note that in principle the RT-DE can be combined with existing real-time MBPT methods and their extension to more advanced self-energy approximations, such as $GW\Gamma$[41].
+This extension is currently underway and will open the doors to the accurate and efficient simulation of time-resolved photo-emission experiments.
 Acknowledgements